# Riccardo Crupi 11 0009-0005-6714-5161    Daniele Regoli 11 $0$-0003-2711-8343    Alessandro Damiano Sabatino 11 $0$-0002-1336-2057    Immacolata Marano 22    Massimiliano Brinis 22    Luca Albertazzi 22    Andrea Cirillo 22    Andrea Claudio Cosentini 11

### v3

@@ Abstract
Explaining outliers occurrence and mechanism of their occurrence can be extremely important in a variety of domains. Malfunctions, frauds, threats, in addition to being correctly identified, oftentimes need a valid explanation in order to effectively perform actionable counteracts. The ever more widespread use of sophisticated Machine Learning approach to identify anomalies make such explanations more challenging. We present the Decision Tree Outlier Regressor (DTOR), a technique for producing rule-based explanations for individual data points by estimating anomaly scores generated by an anomaly detection model. This is accomplished by first applying a Decision Tree Regressor, which computes the estimation score, and then extracting the relative path associated with the data point score. Our results demonstrate the robustness of DTOR even in datasets with a large number of features. Additionally, in contrast to other rule-based approaches, the generated rules are consistently satisfied by the points to be explained. Furthermore, our evaluation metrics indicate comparable performance to Anchors in outlier explanation tasks, with reduced execution time.
@@
Outlier detection Explainability Decision Tree
@@ 1.1 Anomaly detection in the Banking sector - Internal Audit activity
The objectives of internal audit in the banking system are multifaceted and crucial for ensuring the integrity and efficiency of operations. Internal auditors aim to assess and evaluate the effectiveness of internal controls, risk management processes and compliance with regulatory requirements. They also strive to identify and mitigate potential risks, detect fraud, and provide recommendations for improvement [11] [1].
@@
Several anomaly detection techniques have been developed and employed in the banking sector, each offering distinct advantages and trade-offs. Among these techniques, three notable approaches, which leverage diverse mathematical principles, have gained prominence ([17] [7]):
@@ 2 Method
Drawing inspiration from the principles of the Isolation Forest algorithm, our novel XAI method capitalizes on the notion of isolating anomalies with minimal cuts in the feature space. To provide transparent explanations for anomaly detection decisions, we employ decision tree regressors. In our approach, a decision tree regressor is trained to learn the anomaly scores associated with each data point generated by the Anomaly Detector. Notably, we introduce a weighted loss function during training, assigning a significantly higher weight to the data point under scrutiny. This weighting scheme ensures that the decision tree regressor prioritizes accurate estimation of the anomaly score for the target data point, enhancing the interpretability and reliability of the local explanation. Once the decision tree is trained, it is sufficient to extract the path followed by the datapoint to be explained to actually provide an interpretable rule for its anomaly score (algorithm 1).
@@
      /* fit the decision tree to the ‘‘weighted’’ training set */
@@
      rules $\leftarrow extract\_path$(DT, $x_{\text{expl}}$);
@@
In accordance with Ribeiro et al. (2018) [14], we establish notation for a black-box classifier $f$, an instance $x$, and a data distribution $\mathcal{D}$. A condition rule $A_{x}$ comprises a set of feature predicates on $x$ aimed at achieving a level of precision denoted as $\text{precision}(A_{x})$, defined as:
where it is understood that:
$A_{x}(z)=\left\{\begin{aligned} 1&&\text{if}\ z\ \text{satisfies}\ A_{x},\\ 0&&\text{otherwise.}\end{aligned}\right.$ (3)
$\text{validity}(A_{x})=A_{x}(x).$ (4)
@@
The pseudocode for the describred procedure is outlined in algorithm 2.111for an implementation consisting in both generation of the rule and computing the precision refer to the method explain_instances_with_precision in https://github.com/rcrupiISP/DTOR/blob/main/dtor.py.
@@ 3.2 Rule-based XAI
It’s important to note that DTOR estimates the score rather than the binary output of anomaly/non-anomaly. To obtain this output, the same threshold cut employed by the anomaly detection models is provided. While not reported in this work, each rule in the output can provide not only the precision score but also the average anomaly score, which is more informative.
Table 2: Summary of the experiments over seven datasets, three anomaly detectors, and the two compared XAI techniques: Anchors and DTOR. The best performing metric between the two methods is highlighted in bold. Execution time refers to the average time (maximum time over the test set in parenthesis). Precision and coverage are reported with standard deviation over the test set. Validation is reported as a percentage. Rule length is the average length over the explanation rules found for the test set.
@@
banking Execution time 13.17 (19.20) 24.88 (32.24) 17.39 (36.86) 3.79 (5.98) 15.44 (23.75) 14.18 (19.58)
@@
glass Execution time 14.75 (20.94) 3.64 (5.13) 0.13 (0.15) 2.21 (3.51) 0.76 (1.19) 2.49 (3.92)
@@
ionosphere Execution time 169.40 (275.22) 11.25 (13.61) 15.53 (28.24) 3.19 (5.03) 5.76 (8.03) 3.85 (6.62)
@@
lymphography Execution time 16.94 (40.04) 3.90 (5.72) 4.20 (7.98) 2.36 (3.72) 1.07 (1.35) 2.70 (4.40)
@@
musk Execution time 5273.15 (5273.15) 11.92 (16.45) 4.97 (5.19) 11.05 (18.60) 472.87 (1374.55) 15.54 (27.80)
@@
breast cancer Execution time 108.08 (108.08) 4.49 (6.78) 0.50 (0.50) 2.85 (4.65) 4.80 (9.64) 3.67 (6.16)
@@
arrhythmia Execution time - 15.98 (26.03) - 9.70 (17.18) - 13.19 (24.46)

### v4

# Riccardo Crupi and Daniele Regoli and Alessandro Damiano Sabatino and Immacolata Marano and Massimiliano Brinis and Luca Albertazzi and Andrea Cirillo and Andrea Claudio Cosentini 11 0009-0005-6714-5161 11 $0$-0003-2711-8343 11 $0$-0002-1336-2057 2222222211

@@ Abstract
Explaining outliers occurrence and mechanism of their occurrence can be extremely important in a variety of domains. Malfunctions, frauds, threats, in addition to being correctly identified, oftentimes need a valid explanation in order to effectively perform actionable counteracts. The ever more widespread use of sophisticated Machine Learning (ML) approach to identify anomalies make such explanations more challenging. We present the Decision Tree Outlier Regressor (DTOR), a technique for producing rule-based explanations for individual data points by estimating anomaly scores generated by an anomaly detection model. This is accomplished by first applying a decision tree regressor, which computes the estimation score, and then extracting the relative path associated with the data point score. Our results demonstrate the robustness of DTOR even in datasets with a large number of features. Additionally, in contrast to other rule-based approaches, the generated rules are consistently satisfied by the points to be explained. Furthermore, our evaluation metrics indicate comparable performance to Anchors in outlier explanation tasks, with reduced execution time.
@@
Outlier detection, Explainability, Decision Tree
@@ 1.1 Anomaly detection in the Banking sector - Internal Audit activity
The objectives of internal audit in the banking system are multifaceted and crucial for ensuring the integrity and efficiency of operations. Internal auditors aim to assess and evaluate the effectiveness of internal controls, risk management processes and compliance with regulatory requirements. They also strive to identify and mitigate potential risks, detect fraud, and provide recommendations for improvement [11]-[2].
@@
Several anomaly detection techniques have been developed and employed in the banking sector, each offering distinct advantages and trade-offs. Among these techniques, three notable approaches, which leverage diverse mathematical principles, have gained prominence ([17]-[7]):
@@ 2 Method
Drawing inspiration from the principles of the Isolation Forest algorithm, our novel XAI method capitalizes on the notion of isolating anomalies with minimal cuts in the feature space. To provide transparent explanations for anomaly detection decisions, we employ decision tree regressors. In our approach, a decision tree regressor is trained to learn the anomaly scores associated with each data point generated by the Anomaly Detector (AD). Notably, we introduce a weighted loss function during training, assigning a significantly higher weight to the data point under scrutiny. This weighting scheme ensures that the decision tree regressor prioritizes accurate estimation of the anomaly score for the target data point, enhancing the interpretability and reliability of the local explanation. Once the decision tree is trained, it is sufficient to extract the path followed by the datapoint to be explained to actually provide an interpretable rule for its anomaly score (algorithm 1).
In accordance with Ribeiro et al. [14], we establish notation for a black-box classifier $f$, an instance $x$, and a data distribution $\mathcal{D}$ [^1] In this work we can consider this distribution as the empirical distribution of the training set.. A condition rule $A_{x}$ comprises a set of feature predicates on $x$ aimed at achieving a level of precision denoted as $\text{precision}(A_{x})$, defined as:
where it is understood that:
$A_{x}(z)=\left\{\begin{aligned} 1&&\text{if}\ z\ \text{satisfies the rule}\ A_{x},\\ 0&&\text{otherwise.}\end{aligned}\right.$ (3)
For instance, if the rule $A_{x}$ is defined as $A_{x}(z)=(z_{1}>1)\;AND\;(z_{2}>5)$, the candidate sample $\hat{z}=[3,10]$ satisfy the rule, therefore $A_{x}(\hat{z})=1$.
$\text{validity}(A_{x})=A_{x}(x).$ (4)
@@
      /* fit the decision tree to the ‘weighted’ training set */
@@
      rules $\leftarrow$ extract_path(DT, $x_{\text{expl}}$);
@@
The pseudocode for the describred procedure is outlined in algorithm 2.222for an implementation consisting in both generation of the rule and computing the precision refer to the method explain_instances_with_precision() in https://github.com/rcrupiISP/DTOR/blob/main/dtor.py.
@@
To enhance understanding and simplify the process, we present Figure 1, which illustrates the data generation steps.
Figure 1: A simplified illustration of synthetic data generation is presented. Initially, samples from the original dataset are selected based on sub-rules (e.g., $x_{1}>2$ or $x_{2}<0$ in the given example). Subsequently, $N_{\text{gen}}$ samples are drawn for each variable to satisfy the overarching rule $A$. Notably, the image does not depict the discretization of continuous variables or the preservation of inter-variable correlations. However, for illustrative purposes, it is evident that negative values of $x_{3}$ do not occur under rule $A$, as observed in the synthetic dataset.
@@ 3.2 Rule-based XAI
It’s important to note that DTOR estimates the score rather than the binary output of anomaly/non-anomaly. To obtain this output, the same threshold cut employed by the anomaly detection models is provided. While not reported in this work, each rule in the output can provide not only the precision score but also the average anomaly score, which is more informative. Example rule explanations are presented in Table 2, where each rule is associated by its precision, coverage scores and execution time. The comprehensive comparison across multiple datasets and three anomaly detectors is reported in Table 3.
Table 2: Illustrative examples of Isolation Forest explanation on the Lymphography Dataset. The table showcases three dataset samples along with the rules derived from Anchors and DTOR. Each sample is denoted by its corresponding dataset row index, indicating its position within the dataset. While some explanations exhibit similarities, others differ; notably, Anchors often fail to provide any explanation, as observed in example ID 0.
Index row sample Anchors Rule DTOR Rule
36 feature_9 > 1 AND feature_7 > 1 feature_9 > 2 AND feature_7 > 1.5
139 feature_8 <= 1 AND feature_7 > 1 feature_0 > 3.5 AND feature_12 <= 3
0 - feature_4 <= 1.5 AND feature_15 <= 1.5 AND feature_10 > 1.5 AND feature_13 > 3.5
Table 3: Summary of the experiments over seven datasets, three anomaly detectors, and the two compared XAI techniques: Anchors and DTOR. The best performing value between the two methods is highlighted in bold. Execution time refers to the average time (maximum time over the test set in parenthesis). Precision and coverage are reported with standard deviation over the test set. Validation is reported as a percentage. Rule length is the average length over the explanation rules found for the test set.
@@
banking Exec. time 13.17 (19.20) 24.88 (32.24) 17.39 (36.86) 3.79 (5.98) 15.44 (23.75) 14.18 (19.58)
@@
glass Exec. time 14.75 (20.94) 3.64 (5.13) 0.13 (0.15) 2.21 (3.51) 0.76 (1.19) 2.49 (3.92)
@@
ionosphere Exec. time 169.40 (275.22) 11.25 (13.61) 15.53 (28.24) 3.19 (5.03) 5.76 (8.03) 3.85 (6.62)
@@
lymphography Exec. time 16.94 (40.04) 3.90 (5.72) 4.20 (7.98) 2.36 (3.72) 1.07 (1.35) 2.70 (4.40)
@@
musk Exec. time 5273.15 (5273.15) 11.92 (16.45) 4.97 (5.19) 11.05 (18.60) 472.87 (1374.55) 15.54 (27.80)
@@
breast cancer Exec. time 108.08 (108.08) 4.49 (6.78) 0.50 (0.50) 2.85 (4.65) 4.80 (9.64) 3.67 (6.16)
@@
arrhythmia Exec. time - 15.98 (26.03) - 9.70 (17.18) - 13.19 (24.46)